OpenStudy (anonymous):

hello i have a question im almost timmed out!! what is 3 ^2/3 equal to? 3 sqrt9 2sqrt9 3sqrt27 2 sqrt 27

OpenStudy (anonymous):

ill gve medal what every want

OpenStudy (jdoe0001):

\(\bf \large { a^{\frac{{\color{blue} n}}{{\color{red} m}}} \implies \sqrt[{\color{red} m}]{a^{\color{blue} n}} \qquad \qquad \sqrt[{\color{red} m}]{a^{\color{blue} n}}\implies a^{\frac{{\color{blue} n}}{{\color{red} m}}}\qquad thus \\ \quad \\ 3^{\frac{{\color{blue}{ 2 }}}{{\color{red}{ 3 }}}}\implies ? }\)

OpenStudy (anonymous):

..ya what he said

OpenStudy (anonymous):

thankyou u all get fan